# Adaptive Deadline and Batch Layered Synchronized Federated Learning

### v1

@@ Abstract
Federated learning (FL) enables collaborative model training across distributed edge devices while preserving data privacy, and typically operates in a round-based synchronous manner. However, synchronous FL suffers from latency bottlenecks due to device heterogeneity, where slower clients (stragglers) delay or degrade global updates. Prior solutions, such as fixed deadlines, client selection, and layer-wise partial aggregation, alleviate the effect of stragglers, but treat round timing and local workload as static parameters, limiting their effectiveness under strict time constraints. We propose ADEL-FL, a novel framework that jointly optimizes per-round deadlines and user-specific batch sizes for layer-wise aggregation. Our approach formulates a constrained optimization problem minimizing the expected $\ell_{2}$ distance to the global optimum under total training time and global rounds. We provide a convergence analysis under exponential compute models and prove that ADEL-FL yields unbiased updates with bounded variance. Extensive experiments demonstrate that ADEL-FL outperforms alternative methods in both convergence rate and final accuracy under heterogeneous conditions.
1 Introduction
[…]
@@ 1 Introduction
fl is a decentralized machine learning paradigm where multiple clients collaboratively train a shared model under the coordination of a central server, without exchanging their raw data [1]. This approach enables learning from widespread data while preserving privacy and reducing communication by only transmitting model updates. FL has opened opportunities to leverage data from edge devices (smartphones, IoT, etc.) for tasks like text prediction and personalized models, all while keeping sensitive information local [2, 3].
Empirical distributed learning studies showed that some workers are often 5$\times$ slower than the median worker under typical conditions [4]. In synchronous FL (as in the standard federated averaging (FedAvg) protocol [1]), the server must wait for all (or most) clients to send updates before proceeding, so a slow client (called a straggler) can become a bottleneck. This variability in device speed and availability leads to significant latency and can degrade the model if certain clients rarely contribute due to repeated timeouts.
Related Work
@@
A variety of approaches have been proposed to mitigate straggler impact in distributed learning and FL. A simple yet common strategy in synchronous stochastic gradient descent (SGD) is to impose a per-round deadline: the server aggregates whatever updates have arrived by a cutoff time $T$ and ignores late submissions [5]. This drop-stragglers approach has been shown to converge at the same asymptotic rate as full participation [6]. However, discarding stragglers outright can waste useful computation and potentially bias the learned model (since slower devices might consistently be left out). An alternative approach alters federated learning (FL) to be asynchronous, avoiding waiting for stragglers by incorporating each update as it arrives, at the cost of some updates being stale [7]. To reduce the impact of stale updates, staleness-aware aggregation methods were proposed, such as buffering or reweighting updates based on arrival time [8].
@@
Several works aimed at alleviating the harmful effect of stragglers without sacrificing their contributions. Client selection protocols were proposed in [9] to account for expected latency, while potentially reducing the effective data coverage per round. Other methods allow variable workload or model size on different devices by, e.g., training smaller local models on weaker devices [10]. Recently, more refined strategies have leveraged the structure of deep networks by allowing partial model updates. Notably, straggler-aware layer-wise federated learning (SALF)  [11] introduced layer-wise aggregation, where even slow clients contribute gradients for the layers they complete within the deadline. While this improves utilization and convergence, existing methods like SALF typically fix deadlines and batch sizes, missing the potential benefits of dynamically allocating both per-round time and per-client workload based on system heterogeneity.
Research Gap
@@
While prior work addresses straggler mitigation either at the round level (deadlines, client selection) or the local optimization level (layer-wise partial updates), there is no unified approach that optimizes both the deadline and the local optimization to best use the total available time. Yet these decisions are inherently coupled: longer deadlines enable more computation from stragglers, while local optimization procedure affects how much local data each client processes and how long each round takes. Furthermore, existing methods such as SALF default to applying a full FedAvg update when no updates are available for a layer, a practice that can misalign the model under strict time constraints.
Proposed Approach – ADEL-FL
@@
We propose Adaptive Deadline and Batch Layered Federated Learning (ADEL-FL), a novel framework that jointly optimizes per-round deadlines and the mini-batch sizes used for local learning, combined with layer-wise aggregation. In contrast to existing methods, which either fix the deadline per round or treat it as a heuristic hyperparameter, ADEL-FL actively schedules a different deadline for each round, as part of a global optimization objective. Specifically, ADEL-FL formulates the deadline assignment and mini-batch sizing as a constrained optimization problem aimed at minimizing the expected distance to the optimal model under total training time and communication round constraints.
Summary of Contributions:
@@
  • We introduce ADEL-FL, the first FL method to jointly optimize round deadlines and batch sizes with layer-wise aggregation, balancing between reducing stochastic gradient variance and mitigating deadline-induced truncation variance leading to improved convergence under time constraints. Unlike prior methods, our approach accounts for iteration-varying deadlines and modified batch sizes, and it preserves layer parameters when no updates are received.
@@
  • By modeling per-layer computation time using a realistic exponential distribution, which enables analytical estimates of expected updates under deadline constraints and guiding optimal allocation of time and workload, we provide convergence guarantees for ADEL-FL. We rigorously show that its partial updates are unbiased with bounded variance, and analyze its effect on convergence of the learning procedure.
@@
  • We show experimentally that ADEL-FL achieves superior convergence and accuracy compared to existing straggler mitigation approaches under heterogeneous settings. For example, we show an improvement of over 19% in validation accuracy for MNIST [12] and over 4% in CIFAR-10 [13] over existing methods.
@@
The rest of the paper is structured as follows: Section 2 presents the system model and the problem formulation; Section 3 details ADEL-FL algorithm and optimization solution; Section 4 provides experimental results; and Section 5 concludes the paper.
2 Preliminaries and Problem Formulation
@@
In this section, we lay the foundation for deriving ADEL-FL. We begin with modeling synchronous FL in Section 2.1, after which we introduce the system heterogeneity model and discuss its impact on local computation latency in Section 2.2. We then formulate the tackled problem in Section 2.3.
2.1 Federated Learning
@@
In FL [1], a central server trains a model with $n$ parameters denoted ${\bm{w}}\in\mathbb{R}^{n}$, utilizing data stored at a group of $U$ users indexed by $u\in\{1,...,U\}$. Unlike traditional centralized learning, their corresponding datasets-—$\mathcal{D}_{1},\mathcal{D}_{2},...,\mathcal{D}_{U}$—-cannot be transferred to the server due to privacy or communication restrictions [14]. Letting $F_{u}({\bm{w}})$ denote the empirical risk function of the $u$th user, FL aims to find the parameters ${\bm{w}}_{\rm opt}$ which minimize the averaged empirical risk across all users, i.e.,
@@
Broadly speaking, FL follows an iterative procedure operated in rounds. In the $t$th round, the server transmits the current global model ${\bm{w}}_{t}$ to the participating users, who each updates the model using its local dataset and computational resources. These updates are then broadcasted back to the server, which aggregates them to form the new global model; and this process continues until desired convergence is reached. In particular, conventional FL involves training at the devices via local-SGD [15] and server’s aggregation using FedAvg [1]. The former takes the form
${\bm{w}}^{u}_{t+1}\xleftarrow{}{\bm{w}}_{t}-\eta_{t}\nabla F_{u}\left({\bm{w}}_{t};i^{u}_{t}\right),$ (2)
where $i^{u}_{t}$ is the batch index chosen uniformly from all batches of cardinality $S_{u}$ in $\mathcal{D}_{u}$, and $\eta_{t}$ is the learning rate. For FedAvg, the aggregation rule is given by
@@
2.2 Computational Heterogeneity
In FL, participating edge devices often exhibit significant variability in their computational capabilities, resulting in different processing times for computing the local model updates (gradients) in (2) [4]. Formally, considering the $u$th user at global round $t$, let $T_{t}^{u}$ be the computational time required for locally calculating its corresponding gradient $\nabla F_{u}\left({\bm{w}}_{t};i^{u}_{t}\right)$. The individual computational speed $T_{t}^{u}$ can dramatically alter with respect to both $u$ and $t$ due to hardware limitations or additional computational tasks being executed concurrently. Consequently, the overall computational latency per round $t$ in synchronous FL is determined by the slowest device, namely, $\max_{u}T_{t}^{u}$.
This constraint can make synchronous FL abortive for time-sensitive applications, where rapid learning is essential, as discussed in Section 1. This can be tackled, e.g., by imposing a global, per-round, fixed deadline $T_{t}^{d}$ of maximum allowable computation time [5]. Accordingly, nodes not meeting the deadline termed stragglers are dropped in aggregation such that (3) transforms into
${\bm{w}}^{\rm drop}_{t+1}={\bm{w}}^{\rm drop}_{t}-\eta_{t}\frac{1}{|\mathcal{U}_{t}|}\sum_{u\in\mathcal{U}_{t}}\nabla F_{u}({\bm{w}}^{\rm drop}_{t};i^{u}_{t}),\quad\text{where}\quad\mathcal{U}_{t}=\left\{u\mid T_{t}^{u}\leq T_{t}^{d}\right\}.$ (4)
SALF [11] improves upon (4) by, instead of dropping the stragglers, utilizes their partial gradients that are calculated within the deadline limitation, forming the global model in a layer-wise fashion. In particular, for an $L$-layer neural network ${\bm{w}}_{t}=\left[{\bm{w}}_{t}^{1},\dots,{\bm{w}}_{t}^{l},\dots,{\bm{w}}_{t}^{L}\right]^{T}$, (4) changes into updating each layer $l$ separately via
$\tilde{{\bm{w}}}^{l}_{t+1}=\begin{cases}\tilde{{\bm{w}}}^{l}_{t},&\text{if }|\mathcal{U}_{t}^{l}|=0,\quad\text{for}\quad\mathcal{U}_{t}^{l}\subseteq\{1,\dots,U\}\\ \frac{1}{1-p_{l}}\left(\sum\limits_{u\in\mathcal{U}_{t}^{l}}\frac{1}{|\mathcal{U}_{t}^{l}|}\tilde{{\bm{w}}}^{l}_{u,t}-p_{l}\tilde{{\bm{w}}}^{l}_{t}\right)&\text{otherwise},\quad\text{ for}\quad p_{l}\in\mathbb{R},\end{cases}$ (5)
where $p_{l}\triangleq P\left(|\mathcal{U}_{t}^{l}|=0\right)$ is a bias correction constant.
2.3 Problem Formulation
@@
    Total rounds: The FL procedure cannot surpass $R$ global rounds.
@@
    Total training time: The total time required for tuning of the FL model parameters is at most $T_{\max}$. That is, if the execution time of the $t$th FL round is $T_{t}^{d}$, then $\sum_{t=1}^{R}T_{t}^{d}\leq T_{\max}$.
@@
    Devices heterogeneity: Users differ in their computational capabilities, which implies that the latency of computing the gradients for a data batch can greatly vary between users.
3 ADEL-FL
@@
Evidently, Requirements R1-R3 can be satisfied by setting $T_{t}^{d}=T_{\max}/R$, fixing the batch sizes, and employing layer-wise aggregation in (5). Nevertheless, the need to meet R1-R3 while achieving a high-performance FL model learned in dynamically changing environments, motivates an optimized choice of $T_{t}^{d}$ to best leverage the available local resources per round $t$. Moreover, as the local computation time of a user can be altered via its local training hyperparameters, and particularly its batch size, we seek a joint optimization of the per-round deadline and per-user batch size.
The proposed ADEL-FL scheme dynamically combines layer-wise aggregation as in (5) with varying deadline and local optimization, as illustrated in Figure 1. To derive how ADEL-FL sets its deadlines and mini-batch sizes, we first formulate a constrained optimization problem using the expected $\ell_{2}$ distance. Next, in Section 3.2 we derive the bound on the $\ell_{2}$ distance and in Section 3.3 we present the overall ADEL-FL algorithm.
@@
Figure 1: ADEL-FL illustration. Each user performs depth-limited backpropagation, with the extent of computation governed by the deadline. The server aggregates layer-wise gradients across users using only the layers received before the deadline. The color of each layer reflects the set of users that contributed to it, visualized by blending the corresponding user colors. For example, if User $1$ is green and User $U$ is blue, a cyan-colored layer indicates that both contributed to that layer in the current round.
3.1 Optimization Problem

### v2

@@ Abstract
Federated learning (FL) enables collaborative model training across distributed edge devices while preserving data privacy, and typically operates in a round-based synchronous manner. However, synchronous FL suffers from latency bottlenecks due to device heterogeneity, where slower clients (stragglers) delay or degrade global updates. Prior solutions, such as fixed deadlines, client selection, and layer-wise partial aggregation, alleviate the effect of stragglers, but treat round timing and local workload as static parameters, limiting their effectiveness under strict time constraints. We propose Adaptive Deadline and Batch Layered Federated Learning (ADEL-FL), a novel framework that jointly optimizes per-round deadlines and user-specific batch sizes for layer-wise aggregation. Our approach formulates a constrained optimization problem minimizing the expected $\ell_{2}$ distance to the global optimum under total training time and global rounds. We provide a convergence analysis under exponential compute models and prove that ADEL-FL yields unbiased updates with bounded variance. Extensive experiments demonstrate that ADEL-FL outperforms alternative methods in both convergence rate and final accuracy under heterogeneous conditions.
I Introduction
@@
Federated learning (FL) is a decentralized machine learning paradigm where multiple clients collaboratively train a shared model under the coordination of a central server, without exchanging their raw data [1]. This approach enables learning from widespread data while preserving privacy and reducing communication by only transmitting model updates. FL has opened opportunities to leverage data from edge devices (smartphones, IoT, etc.) for training models while keeping sensitive information local [2, 3].
FL typically operates as a round based iterative procedure [4]. In each round, the central server transmits the current global model to a set of participating clients. Each client then uses its local dataset and computational resources to compute an update to the model via, e.g., stochastic gradient descent (SGD) [5]. Following local computation, these model updates are transmitted back to the server, which aggregates them to form the new global model. In the standard federated averaging (FedAvg) protocol [1], the server must wait for all (or most) clients to send updates before proceeding, so a slow client (called a straggler) can become a bottleneck. This operation combined with the variability in device speed and availability can lead to significant latency and degrade the model if certain clients rarely contribute due to repeated timeouts.
In practical federated systems, client delays in training arise from two main sources: $(i)$ the time clients spend computing local updates; and $(ii)$ the time required to communicate these updates to the server [6, 7]. The communication delay can be partially mitigated by techniques such as sparsification [8, 9, 10, 11] and compression [12, 13, 14, 15] of the model updates before transmission. The computational delay often remains a dominant bottleneck in synchronous FL, as empirical distributed learning studies showed that some workers are often significantly slower than the median worker under typical conditions [16].
@@
A variety of approaches have been proposed to mitigate straggler impact in distributed learning and FL. A simple yet common strategy in synchronous SGD is to impose a per-round deadline: the server aggregates whatever updates have arrived by a cutoff time $T$ and ignores late submissions [17]. This drop-stragglers approach has been shown to converge at the same asymptotic rate as full participation [18]. However, discarding stragglers outright can waste useful computation and potentially bias the learned model (since slower devices might consistently be left out). An alternative approach alters FL to be asynchronous [19], avoiding waiting for stragglers by incorporating each update as it arrives, at the cost of some updates being stale [20]. Stale updates, i.e., gradients computed on an older version of the global model, can introduce gradient mismatch and slow down or even destabilize convergence [21]. To reduce the impact of stale updates, staleness-aware aggregation methods were proposed, such as buffering or reweighting updates based on arrival time [22]. Still, asynchronous FL typically requires the number of slow users to be small for stable learning [23], and is thus limited when learning under tight latency constraints. It was proposed in [24, 25] to combine asynchronous FL with synchronous operation by synchronizing local stale models while performing global updating asynchronously. However, such semi-asynchronous FL, there may still be high variation between participating users in their local computation times, which again limits operation under tight latency constraints.
@@
Several works aimed at alleviating the harmful effect of stragglers without sacrificing their contributions. Client selection protocols were proposed to account for expected latency [26], share calculation time between users and server [27], and assign lesser weight to older updates [28], while potentially reducing the effective data coverage per round and risking systemic bias if the same set of low-latency clients is repeatedly selected [29]. Other methods allow variable workload or model size on different devices by, e.g., training smaller local models on weaker devices [30]. Recently, more refined strategies geared towards tight latency FL have leveraged the structure of deep networks by allowing partial model updates. Notably, straggler-aware layer-wise federated learning (SALF) [31] introduced layer-wise aggregation, where even slow clients contribute gradients for the layers they complete within the deadline. This layer-wise approach is well-suited for deep neural networks where backpropagation is inherently sequential. While this improves utilization and convergence, existing methods like SALF typically assumed fixed deadlines and batch sizes, missing the potential benefits of dynamically allocating both per-round time and per-client workload based on system heterogeneity.
@@
While prior work addresses straggler mitigation either at the round level (deadlines, client selection) or the local optimization level (layer-wise partial updates), there is no unified approach that optimizes both the deadline and the local optimization to best use the total available time. Yet, these decisions are inherently coupled: longer deadlines enable more computation from stragglers, while local optimization procedure affects how much local data each client processes and how long each round takes. Furthermore, existing methods such as SALF default to applying a full FedAvg update when no updates are available for a layer, a practice that can misalign the model under strict time constraints.
@@
We propose Adaptive Deadline and Batch Layered Federated Learning (ADEL-FL), a novel framework that jointly optimizes per-round deadlines in synchronous FL and the mini-batch sizes used for local learning, combined with layer-wise aggregation. In contrast to existing methods, which either fix the deadline per round [17] or treat it as a heuristic hyperparameter [32], ADEL-FL actively schedules a different deadline for each round, as part of a global optimization objective. Specifically, ADEL-FL formulates the deadline assignment and mini-batch sizing as a constrained optimization problem aimed at minimizing the expected distance to the optimal model under total training time and communication round constraints.
Summary of Contributions
@@
  • Optimized Straggler-Adaptive Synchronous FL Algorithm: We introduce ADEL-FL, the first FL method to jointly optimize round deadlines and batch sizes with layer-wise aggregation, balancing between reducing stochastic gradient variance and mitigating deadline-induced truncation variance leading to improved convergence under time constraints. Unlike prior methods, our approach accounts for iteration-varying deadlines and modified batch sizes, and it preserves layer parameters when no updates are received.
@@
  • Theoretical Analysis: By modeling per-layer computation time using a realistic exponential distribution, which enables analytical estimates of expected updates under deadline constraints and guiding optimal allocation of time and workload, we provide convergence guarantees for ADEL-FL. We rigorously show that its partial updates are unbiased with bounded variance, and analyze its effect on convergence of the learning procedure.
@@
  • Extensive Experimentation: We show experimentally that ADEL-FL achieves superior convergence and accuracy compared to existing straggler mitigation approaches under heterogeneous settings. For example, we show an improvement of over 19% in validation accuracy for MNIST [33] and over 13% in CIFAR-10 [34] over existing methods.
@@
The rest of the paper is structured as follows: Section II presents the system model and the problem formulation; Section III details ADEL-FL; Section IV provides experimental results; and Section V concludes the paper.
II Preliminaries and Problem Formulation
@@
In this section, we lay the foundation for deriving ADEL-FL. We begin with modeling synchronous FL in Section II-A, after which we introduce the system heterogeneity model and discuss its impact on local computation and communication latency in Section II-B. We then formulate the tackled problem in Section II-C.
II-A Federated Learning
@@
In FL [1], a central server trains a model with $n$ parameters denoted ${\bm{w}}\in\mathbb{R}^{n}$, utilizing data stored at a group of $U$ users indexed by $u\in\{1,...,U\}$. Unlike traditional centralized learning, their corresponding datasets—denoted $\mathcal{D}_{1},\mathcal{D}_{2},...,\mathcal{D}_{U}$—cannot be transferred to the server due to privacy or communication restrictions [4]. Letting $F_{u}({\bm{w}})$ denote the empirical risk function of the $u$th user, FL aims to find the parameters ${\bm{w}}_{\rm opt}$ which minimize the averaged empirical risk across all users, i.e.,
@@
Broadly speaking, FL follows an iterative procedure operated in rounds. In the $t$th round, the server transmits the current global model ${\bm{w}}_{t}$ to the participating users, who each updates the model using its local dataset and computational resources. These updates are then broadcasted back to the server, which aggregates them to form the new global model; and this process continues until desired convergence is reached. In particular, conventional FL involves training at the devices via local SGD [5] and server’s aggregation using FedAvg [1]. The former takes the form
${\bm{w}}^{u}_{t+1}\xleftarrow{}{\bm{w}}_{t}-\eta_{t}\nabla F_{u}\left({\bm{w}}_{t};i_{t}^{u}\right),$ (2)
where $i_{t}^{u}$ is the batch index chosen uniformly from all batches of cardinality $S_{t}^{u}$ in $\mathcal{D}_{u}$, and $\eta_{t}$ is the learning rate. For FedAvg, the aggregation rule is given by
@@
II-B Computation and Communication Heterogeneity
In FL, participating edge devices often exhibit significant variability in their computational and communication capabilities, resulting in different processing times for computing the local model updates (gradients) in (2) and uploading them to the server [6, 7]. Formally, considering the $u$th user at global round $t$, let $T_{t}^{{\rm b},u}$ be the computational time required for backpropagation and locally calculating its corresponding gradient $\nabla F_{u}\left({\bm{w}}_{t};i_{t}^{u}\right)$. The individual computational speed $T_{t}^{{\rm b},u}$ can dramatically alter with respect to both $u$ and $t$ due to hardware limitations or additional computational tasks being executed concurrently. In addition, after completing the local computation, each device must transmit its updated model to the server in communication time denoted by $T_{t}^{{\rm c},u}$ at round $t$ for user $u$. The total iteration time $T_{t}^{u}$ of user $u$ is given by: $T_{t}^{u}=T_{t}^{{\rm b},u}+T_{t}^{{\rm c},u}$. Consequently, the overall computational latency per round $t$ in synchronous FL is determined by the slowest device, namely, $\max_{u}T_{t}^{u}$. This constraint can make synchronous FL abortive for time-sensitive applications, where rapid learning is essential. This can be tackled, e.g., by imposing a global, per-round, fixed deadline $T_{t}^{\rm d}$ of maximum allowable computation time [17]. Accordingly, nodes not meeting the deadline termed stragglers are dropped in aggregation such that (3) transforms into
${\bm{w}}^{\rm drop}_{t+1}$ $={\bm{w}}^{\rm drop}_{t}-\eta_{t}\frac{1}{|\mathcal{U}_{t}|}\sum_{u\in\mathcal{U}_{t}}\nabla F_{u}({\bm{w}}^{\rm drop}_{t};i_{t}^{u}),$ (4)
where $\quad\mathcal{U}_{t}=\left\{u\mid T_{t}^{u}\leq T_{t}^{\rm d}\right\}.$
SALF [31] improves upon (4) by, instead of dropping the stragglers, utilizes their partial gradients that are calculated within the deadline limitation, forming the global model in a layer-wise fashion. In particular, for an $L$-layer neural network ${\bm{w}}_{t}=\left[{\bm{w}}_{t}^{1},\dots,{\bm{w}}_{t}^{l},\dots,{\bm{w}}_{t}^{L}\right]^{T}$, and user updates ${\bm{w}}_{u,t}^{l}$ (4) changes into updating each layer $l$ separately via
$\tilde{{\bm{w}}}^{l}_{t+1}=\begin{cases}\tilde{{\bm{w}}}^{l}_{t},&\text{if }|\mathcal{U}_{t}^{l}|=0,\quad\\ \frac{1}{1-p_{t}^{l}}\left(\sum\limits_{u\in\mathcal{U}_{t}^{l}}\frac{1}{|\mathcal{U}_{t}^{l}|}{\bm{w}}^{l}_{u,t}-p_{t}^{l}\tilde{{\bm{w}}}^{l}_{t}\right)&\text{otherwise},\end{cases}$ (5)
where $p_{t}^{l}\triangleq P\left(|\mathcal{U}_{t}^{l}|=0\right)$ is a bias correction constant.
II-C Problem Formulation
@@
    Total rounds: The FL procedure cannot surpass a predefined number $R$ of global rounds.
@@
    Total training time: The total time required for learning the model parameters is at most $T_{\max}$. That is, if the execution time of the $t$th FL round is $T_{t}^{\rm d}$, then $\sum_{t=1}^{R}T_{t}^{\rm d}\leq T_{\max}$.
@@
    Devices heterogeneity: Users differ in their computational capabilities, thus the latency of computing the gradients for a data batch can greatly vary between users.
III ADEL-FL
@@
Evidently, Requirements R1-R3 can be satisfied by setting $T_{t}^{\rm d}=T_{\max}/R$, fixing the batch sizes, and employing layer-wise aggregation in (5). Nevertheless, the need to meet R1-R3 while achieving a high-performance FL model learned in dynamically changing environments, motivates an optimized choice of $T_{t}^{\rm d}$ to best leverage the available local resources per round $t$. Moreover, as the local computation time of a user can be altered via its local training hyperparameters, and particularly its batch size, in this work, we seek a joint optimization of the per-round deadline and per-user batch size. The proposed ADEL-FL scheme dynamically combines layer-wise aggregation as in (5) with varying deadline and local optimization, as illustrated in Figure 1. To derive how ADEL-FL sets its deadlines and mini-batch sizes, in Section III-A we first formulate a constrained optimization problem using the expected $\ell_{2}$ distance. Next, in Section III-B we derive the bound on the $\ell_{2}$ distance and in Section III-C we present the overall ADEL-FL algorithm.
@@
Figure 1: ADEL-FL illustration. Initially, the server executes an optimization step to determine the adaptive per-round deadlines $\{T_{t}^{\rm{d}}\}_{t=1}^{R}$, represented by the varying amount of sand in the hourglass. Then, for every round, each user performs depth-limited backpropagation, with the extent of computation governed by the deadline. The server aggregates layer-wise gradients across users using only the layers received before the deadline. The color of each layer reflects the set of users that contributed to it, visualized by blending the corresponding user colors. For example, if User 11 is green and User $U$ is blue, a cyan-colored layer indicates that both contributed to that layer in the current round.
III-A Optimization Problem